OpenStudy (anonymous):

if x varies directly as y, and x = 48 when y = 8, find x when y = 10

OpenStudy (tamtoan):

varies directly mean...as x increase, y increase as well...and normally the relation you can write as y = c.x where c is a constant ...from here, with those values you have there, can you figure out?

OpenStudy (anonymous):

im clueless

OpenStudy (tamtoan):

y = c.x when x = 48, y = 8 ...can you figure out c from here?

OpenStudy (anonymous):

no im sorry i cant.

OpenStudy (anonymous):

c*8=48 c*10=x

OpenStudy (anonymous):

c= 6

OpenStudy (anonymous):

6*8=48 6*10=60 so x=60 when y=10

OpenStudy (anonymous):

thank you both.

OpenStudy (anonymous):

your welcome